Two bodies of masses 10,000kg and 1000kg respectively are seperated by 1m. At what distance from the smaller body. the intenstisity of the gravitational field will be zero

Grade:11

2 Answers

Chetan Mandayam Nayakar
11 years ago

let answer=r

1000/r2=10000/(1-r)2

orange panda
11 years ago

GM1/D^2=GM2/(1-D)^2.  solve for D